In a coffee shop there are 10 people sharing an access point with a rate speed of 2Gbps,The throughput is ½ the rate speed.Several people are downloading . Each is getting an average of 100Mbps. How many people are using the Internet at that moment.
Answer #1

Throughput =( 1/2)* rate speed

= (1/2)*2 Gbps

= 1Gbps

Each internet user is downloading data at rate of 100 Mbps

So, say x number of people are using internet then

x * 100 Mbps = 1Gbps

x = 109 bps / (100 * 106 bps)

x = 109-8

x = 10

So,the number of people using internet at current moment = x = 10